Using polar coordinates find if the limit does or does not exist for


Lim(x,y)approaches(0,0) (2x/(x^2+X+y^2))

1
Expert's answer
2022-05-20T08:30:25-0400

Let "x=r\cos \theta, y=r\sin\theta." Then


"\dfrac{2x}{x^2+x+y^2}=\dfrac{2r\cos \theta}{r^2+r\cos \theta}=\dfrac{2\cos \theta}{r+\cos \theta}"

The limit exists and


"\lim\limits_{(x,y)\to(0,0)}\dfrac{2x}{x^2+x+y^2}=\lim\limits_{r\to0}\dfrac{2\cos \theta}{r+\cos \theta}"

"=\dfrac{2\cos \theta}{0+\cos \theta}=2"